OpenStudy (anonymous):

Please help! I will medal! Thanx! :) Which is the equation of a circle with center (-2,3) and radius r=5? a. (x+2)^2 + (y-3)^2= 10 b. (x+2)^2 + (y-3)^2= 25 c. (x-2)^2 + (y+3)^2= 10 d. (x-2)^2 + (y+3)^2=25

OpenStudy (anonymous):

@raiders88

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

oh thanks!

Nnesha (nnesha):

\[\huge\rm (x-h)^2 + (y-k)^2 = r^2\] equation of circle in standard form where ( h,k) is the center and r is radius just replace variables by their values

OpenStudy (anonymous):

so would i put (x+2)^2 = (y-3^)2= for the equation?how would i get the ending answer?

Nnesha (nnesha):

r=5 what is r^2 = ?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

so it would be B then?

OpenStudy (anonymous):

thank you!

Nnesha (nnesha):

great !! :-)

Nnesha (nnesha):

my pleasure :-)

OpenStudy (anonymous):

could you help me with a few more?

Nnesha (nnesha):

i'll try :-)

OpenStudy (anonymous):

thx :) i'll post the next one on here hang on :)

Nnesha (nnesha):

okay

OpenStudy (anonymous):

A circle with center (-1,2) passes through point (2,-2). Which is true? a. the radius is sqrt5 b. the diameter is 10 c. the equation is (x+1)^2 + (y-2)^2= 10 d. the circumference is 25pi

Nnesha (nnesha):

okay now again circle is (h,k) points passing through (x , y) so replace x and y by (2,-2) and (h,k) by (-1,2) then solve for r

Nnesha (nnesha):

same equation \[\huge\rm (x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (anonymous):

how would i solve for r after i get the equation?

OpenStudy (anonymous):

so i got (2-(-1)^2 + (-2 - 2)^2 then i solved and got 9 + 16 = 25

Nnesha (nnesha):

same equation \[\huge\rm (\color{green}{x}-\color{blue}{h})^2 + (\color{red}{y}-\color{gold}{k})^2 = r^2\] \[\huge\rm (\color{green}{2},\color{red}{-2})\] \[\huge\rm (\color{blue}{-1},\color{gold}{2})\] plug in x,y,h,k values

OpenStudy (anonymous):

so would the answer be either a or d right?

OpenStudy (anonymous):

i think the answer is a

Nnesha (nnesha):

ohhh

OpenStudy (anonymous):

but then again the square root of 5 is 2.24 so that isn't correct.

Nnesha (nnesha):

r^2 = 16+9 r^2 = 25 to solve for r you should take square root both side so \[\huge\rm \sqrt{r^2}= \sqrt{25}\] r= ?

OpenStudy (anonymous):

the square root of 5 is correct then!

OpenStudy (anonymous):

right?

Nnesha (nnesha):

 25 is perfect square root which is equal to 5

Nnesha (nnesha):

 hint :radius is half of diameter